19.62 per mach. hour Now calculate the total overhead allocated to each job under a simple costing system. (Round your answer to the nearest cent) Simple costing system Job 220 Job 330 Indirect cost allocaterd 588.60 S 1,373.40 Requirement 2. Compute the total overhead allocated to each job under an activity-based costing system using the appropriate activity drivers (Rou the allocation rates to three decimal places, X.X0X. Abbreviations used:bgt budgoted, mach. machine, maint.. purch. purchase, aty. quantity, prod. produced.) maintonanoe, ma! - material ABC allocation rate Choose from any drop-down list and then click Check Answer Clear All Chack Answer (A
media%2F87f%2F87fd5f2f-9708-48c4-950f-0c
media%2F1c6%2F1c66868f-543e-4bfc-a0c7-b9
media%2F4e8%2F4e880fe9-c313-4d08-bd35-74
Answer #1

ABC allocation Rate = Budgeted cost of activity /Budgeted total qty of cost allocation base.

Overhead allocation
Activity Activity rates Job 220 Job 330
Purchasing 28500/1500=19 per purchase order 19*21=399 19*9=171
Material handling 47700/5300=9 per material moves 9*18=162 9*6=54
machine maintenance 100000/10000=10 per machine hours 30*10=300 70*10=700
product inspection 6800/1700=4 per inspections 10*4=40 4*2=8
packaging 13200/3300=4 per unit produced 4*17=68 4*5=20
Overhead allocated 969 953